OpenStudy (anonymous):

Perfom indicated operation and, if possible, simplify. Please show all work. -2/3-(-2/7)

OpenStudy (lilg132):

-8/21

OpenStudy (lilg132):

-2/3- (-2/7) -0.6666666667 + 0.2857142857 = -0.380952381 -2/3 + 2/7 = -8/21

OpenStudy (lilg132):